# algorithmic

### v1

@@ Abstract
We propose a unifying framework for methods that perform Bayesian online learning in non-stationary environments. We call the framework BONE, which stands for (B)ayesian (O)nline learning in (N)on-stationary (E)nvironments. BONE provides a common structure to tackle a variety of problems, including online continual learning, prequential forecasting, and contextual bandits. The framework requires specifying three modelling choices: (i) a model for measurements (e.g., a neural network), (ii) an auxiliary process to model non-stationarity (e.g., the time since the last changepoint), and (iii) a conditional prior over model parameters (e.g., a multivariate Gaussian). The framework also requires two algorithmic choices, which we use to carry out approximate inference under this framework: (i) an algorithm to estimate beliefs (posterior distribution) about the model parameters given the auxiliary variable, and (ii) an algorithm to estimate beliefs about the auxiliary variable. We show how this modularity allows us to write many different existing methods as instances of BONE; we also use this framework to propose a new method. We then experimentally compare existing methods with our proposed new method on several datasets; we provide insights into the situations that make one method more suitable than another for a given task.
@@ 1 Introduction
Our goal in this paper is to study adaptive probabilistic methods that learn to make accurate predictions about the next output ${\bm{y}}_{t+1}$ based on the next input ${\bm{x}}_{t+1}$ and a sequence of past inputs and outputs, $({\bm{x}}_{1:t},{\bm{y}}_{1:t})$, where $t$ indexes time. This is often called prequential forecasting (Gama et al., 2008).
@@
Many prediction methods assume that the data generating process (DGP) $p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},{\bm{x}}_{1:t},{\bm{y}}_{1:t})$ is static through time. However, real-world data often comes from non-stationary distributions, where the underlying data distribution changes, either gradually (e.g., rising mean temperature) or abruptly (e.g., price shocks after a major news event). In this paper we propose a unified framework for tackling (one-step-ahead) forecasts in such (potentially) non-stationary environments. Our framework encompasses many existing lines of work, including online continual learning (Dohare et al., 2024), prequential forecasting (Liu, 2023), test-time adaptation (Schirmer et al., 2024), and neural contextual bandits (Riquelme et al., 2018).
The framework we propose in this paper, which we call BONE — which stands for (B)ayesian (O)nline learning in (N)on-stationary (E)nvironments — is based on a form of Bayesian inference in a hierarchical model,111 It is worth clarifying what we mean by “a form of Bayesian inference”. We focus on adaptive rules, which are defined as “accumulating experience about the properties of the environment and leveraging this experience to improve the model’s performance” (Peterka, 1981). While the adaptive rules that we consider in this work align with Bayes’ rule, this alignment does not imply that we adopt the full set of assumptions required to be strictly Bayesian. By strictly Bayesian, we mean a framework in which the properties of the data-generating process are fully known, well specified, and described within a formal mathematical model with uncertainty quantified about all unknowns (Faden & Rausser, 1976; Peterka, 1981; Knoblauch et al., 2022). Thus, our approach uses Bayes’ rule only to provide a rational basis for sequential decision-making while preserving adaptability. and is composed of three modelling choices and two algorithmic choices. The modelling choices are: (M.1) a model for the measurements, (M.2) an auxiliary process to model non-stationarity, and (M.3) a prior over model parameters conditioned on the auxiliary process and the past data. The algorithmic choices are: (A.1) an algorithm to compute an approximate posterior over the model parameters given the auxiliary variables, and (A.2) an algorithm to compute an approximate posterior —or more generally, a set of weights— over the auxiliary variables. We show how these different axes of variation span a wide variety of existing and new methods (see Table 2). We then perform an experimental comparison on a variety of tasks. Specifically, we consider prequential forecasting, classification, bandits, and unsupervised time-series segmentation. An easy-to-use library that implements these methods, written in Jax (Bradbury et al., 2018), is available at https://github.com/gerdm/BONE.
@@ 2 The framework
2.1 The stationary case
@@
Consider a sequence of target measurements ${\bm{y}}_{1:t}=({\bm{y}}_{1},\ldots,{\bm{y}}_{t})$ with ${\bm{y}}_{i}\in^{d}$, and features ${\bm{x}}_{1:t}=({\bm{x}}_{1},\ldots,{\bm{x}}_{t})$ with ${\bm{x}}_{i}\in^{q}$. Let $\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$ be a datapoint and $\mathcal{D}_{1:t}=(\mathcal{D}_{1},\ldots,\mathcal{D}_{t})$ the dataset at time $t$. We are interested in methods that efficiently compute $p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},\mathcal{D}_{1:t})$ in a recursive, online way. In our setting, one observes ${\bm{x}}_{t+1}$ just before observing ${\bm{y}}_{t+1}$; thus, to make a prediction about ${\bm{y}}_{t+1}$, we have ${\bm{x}}_{1:t+1}$ and ${\bm{y}}_{1:t}$ at our disposal.222 The input features ${\bm{x}}_{t+1}$ and output measurements ${\bm{y}}_{t+1}$ can correspond to different time steps. For example, ${\bm{x}}_{t+1}$ can be the state of the stock market at a fixed date and ${\bm{y}}_{t+1}$ is the return on a stock some days into the future.
Our estimate of the mean of $p({\bm{y}}_{t+1}|{\bm{x}}_{t+1})$ is encoded in a parametric model $h({\bm{\theta}}_{t};\,{\bm{x}}_{t+1})$ (e.g., a neural network), where ${\bm{\theta}}_{t}\in^{m}$ is the parameter vector. Since the parameters are unknown, we adopt a Bayesian approach and write the prediction function as
$\mathbb{E}[h({\bm{\theta}}_{t};\,{\bm{x}}_{t+1})\,|\,\mathcal{D}_{1:t}]=\int h({\bm{\theta}}_{t};\,{\bm{x}}_{t+1})\,\lambda({\bm{\theta}}_{t};\,\mathcal{D}_{1:t})\mathrm{d}{\bm{\theta}}_{t},$ (1)
where $\lambda({\bm{\theta}}_{t};\,\mathcal{D}_{1:t})$ is a density over the model parameters ${\bm{\theta}}_{t}\in^{m}$. In what follows we take $\lambda({\bm{\theta}}_{t};\,\mathcal{D}_{1:t})$ to be the Bayesian posterior density (or an approximation)
$\lambda({\bm{\theta}}_{t};\,\mathcal{D}_{1:t})=p({\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t}).$ (2)
Next, we modify the above static framework to account for non-stationarity.
2.2 BONE: the non-stationary case
The class of methods described above works well when the data-generating process is well-specified. In practice, however, this is often not the case. Thus, to adapt to regime changes and other forms of non-stationarity, we introduce an auxiliary random variable $\psi_{t}\in\boldsymbol{\Psi}_{t}$, where $\boldsymbol{\Psi}_{t}$ is a set of possible values of the auxiliary variable $\psi_{t}$. The purpose of $\psi_{t}$ is to encode information about the current regime so this can be used to predict the upcoming measurement. This can be achieved, for instance, by determining which past data points most closely align with the most recent measurements or by selecting a model —from a set of candidate models— that predicts the future observation. We describe the auxiliary variable in detail in Section 2.6.
Given $\mathcal{D}_{1:t}$, ${\bm{x}}_{t+1}$, and $\psi_{t}$, a prediction for ${\bm{y}}_{t+1}$ is
$\hat{{\bm{y}}}_{t+1}^{(\psi_{t})}=\mathbb{E}_{\lambda_{t}}[h({\bm{\theta}}_{t};\,{\bm{x}}_{t+1})\,|\,\psi_{t}]:=\int h({\bm{\theta}}_{t};\,{\bm{x}}_{t+1})\,\lambda({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})\mathrm{d}{\bm{\theta}}_{t}\,.$ (3)
@@ 2.2 BONE: the non-stationary case
Here, we use the shorthand notation $\lambda_{t}$ for $\lambda({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})$, and we write $\mathbb{E}_{\lambda_{t}}[\cdot\,|\,\psi_{t}]$ and $\hat{{\bm{y}}}_{t+1}^{(\psi_{t})}$ to draw attention to the dependence on $\psi_{t}$. Given that $\lambda_{t}$ is the Bayesian posterior,
$\lambda({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})\propto\tau({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t-1})\,p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t}),$ (4)
where $\tau$ is the conditional prior density. This is a modelling choice that imposes an inductive bias on the model parameters based on the data $\mathcal{D}_{1:t-1}$ and the value $\psi_{t}$. The term $p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})$ is the likelihood function, which is assumed to have the property that for any $t$
$\mathbb{E}\left[{\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t}\right]=\int{\bm{y}}_{t}\,p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})\,\mathrm{d}{\bm{y}}_{t}=h({\bm{\theta}}_{t};{\bm{x}}_{t}).$ (5)
Finally, we introduce the function $\nu_{t}(\psi_{t})$ that weights the importance of each $\psi_{t}\in\boldsymbol{\Psi}_{t}$ when making an estimate of ${\bm{y}}_{t+1}$. More precisely, we use $\nu_{t}(\psi_{t})$ to make the estimation $\hat{{\bm{y}}}_{t+1}$ as follows
$\hat{{\bm{y}}}_{t+1}$ $=\int\left(\int h({\bm{\theta}}_{t};\,{\bm{x}}_{t+1})\,\lambda_{t}({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})\,\mathrm{d}{\bm{\theta}}_{t}\right)\,\nu_{t}(\psi_{t})\,\mathrm{d}\psi_{t}$ (6)
$=:\mathbb{E}_{\nu_{t}}\Big{[}\,\mathbb{E}_{\lambda_{t}}[h({\bm{\theta}}_{t};\,{\bm{x}}_{t+1})\,|\,\psi_{t}]\,\Big{]}\,.$
In the Bayesian approach, we use $\nu_{t}(\psi_{t})=p(\psi_{t}|\mathcal{D}_{1:t})$, but we also consider other ways of defining this measure weighting term in order to accommodate non-Bayesian approaches in the literature, as we discuss in Section 2.6.
As we show in this paper, many methods in the literature can be written as instances of (6). In particular, these methods depend on the modelling choices of: (M.1: likelihood) the model for the measurements in the form of the density $p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})$ which in turns determines $h$ through (5); (M.2: auxvar) an auxiliary variable $\psi_{t}$ that modulates the predictions when non-stationarity is present; (M.3: prior) a conditional prior density $\tau$ over model parameters conditioned on past information and the auxiliary variable $\psi$. Having specified (M.1 – M.3) inference of the model is given by (A.1: posterior) a method to compute or approximate $\lambda_{t}$ in (4), and (A.2: weighting) a method to compute or approximate $\nu_{t}$ which weights possible values of $\psi_{t}$.
@@
Algorithm 1 presents pseudocode for the prediction and update steps of a method within the BONE framework. Notably, these components can be broadly divided into two categories: modeling and algorithmic. The modelling components determine the inductive biases in the model, and correspond to $h$, $\psi_{t}$, and $\tau$. The algorithmic components dictate how operations are carried out to produce a final prediction — this corresponds to $\lambda_{t}$ and $\nu_{t}$.
@@
0:  $p({\bm{y}}\,|\,{\bm{\theta}},{\bm{x}}_{t})$ // Choice of (M.1: likelihood)
@@
2:     $\tau_{t}({\bm{\theta}}_{t};\,\psi_{t})\leftarrow\tau({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t-1})$ // choice of (M.3: prior)
3:     $\lambda_{t}({\bm{\theta}}_{t};\,\psi_{t})\leftarrow\lambda({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t})\propto\tau_{t}({\bm{\theta}}_{t}\;\psi_{t})\,p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})$// choice of (A.1: posterior)
4:     $\nu_{t}(\psi_{t})\leftarrow\nu(\psi_{t}\,|\,\mathcal{D}_{1:t})$ // choice of (A.2: weighting)
5:     $\hat{{\bm{y}}}_{t+1}^{(\psi_{t})}\leftarrow\mathbb{E}_{\lambda_{t}}[h({\bm{\theta}}_{t};\,{\bm{x}}_{t+1})\,|\,\psi_{t}]$ // conditional prequential prediction

### v3

@@ Abstract
We propose a unifying framework for methods that perform probabilistic online learning in non-stationary environments. We call the framework BONE, which stands for generalised (B)ayesian (O)nline learning in (N)on-stationary (E)nvironments. BONE provides a common structure to tackle a variety of problems, including online continual learning, prequential forecasting, and contextual bandits. The framework requires specifying three modelling choices: (i) a model for measurements (e.g., a neural network), (ii) an auxiliary process to model non-stationarity (e.g., the time since the last changepoint), and (iii) a conditional prior over model parameters (e.g., a multivariate Gaussian). The framework also requires two algorithmic choices, which we use to carry out approximate inference under this framework: (i) an algorithm to estimate beliefs (posterior distribution) about the model parameters given the auxiliary variable, and (ii) an algorithm to estimate beliefs about the auxiliary variable. We show how the modularity of our framework allows for many existing methods to be reinterpreted as instances of BONE, and it allows us to propose new methods. We compare experimentally existing methods with our proposed new method on several datasets, providing insights into the situations that make each method more suitable for a specific task. We provide a Jax open source library to facilitate the adoption of this framework.
@@ 1 Introduction
In this paper, we unify adaptive probabilistic methods that learn to make predictions about the next output ${\bm{y}}_{t+1}$ based on the next input ${\bm{x}}_{t+1}$ and a sequence of past inputs and outputs, $({\bm{x}}_{1:t},{\bm{y}}_{1:t})$, where $t$ indexes time. This is often called prequential forecasting (Gama et al., 2008), online learning (Zhang, 2023, Chapter 1), or sequential online prediction (Liu, 2023).
@@
Many probabilistic prediction methods assume that the data generating process (DGP) $p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},{\bm{x}}_{1:t},{\bm{y}}_{1:t})$ is static through time. However, real-world data often comes from non-stationary distributions, where the underlying data distribution changes, either gradually (e.g., rising mean temperature) or abruptly (e.g., price shocks after a major news event). In this paper we propose a unified framework for tackling (one-step-ahead) forecasts in such (potentially) non-stationary environments. Building on recent advances in generalised (pseudo) Bayesian methods (Bissiri et al., 2016; Knoblauch et al., 2022; Khan & Rue, 2023), we show that our framework unifies classical Bayesian approaches and naturally extends to methods derived from Bayes’ rule. We demonstrate that all these methods are related to a three-layered hierarchical state-space model, where the top layer captures changes in the data, the second layer governs the evolution of model parameters, and the third layer represents the observations.
Our framework accommodates and extends many existing lines of work, including online continual learning (Dohare et al., 2024), prequential (one-step-ahead) forecasting (Liu, 2023), test-time adaptation (Schirmer et al., 2024), neural contextual bandits (Riquelme et al., 2018), filtering (Basseville et al., 1993), and changepoint detection or segmentation (Gupta et al., 2024). For example, methods that were originally developed to tackle changepoint detection can be reinterpreted and applied within our framework to address a wider range of sequential online learning problems. This flexibility allows us to gather tools that were designed for specific tasks—often studied in isolation within distinct subfields—into a unifying framework for handling non-stationarity across various application domains.
The framework we propose in this paper, which we call BONE—which stands for (B)ayesian (O)nline learning in (N)on-stationary (E)nvironments—is based on a form of generalised Bayesian inference in a hierarchical model, and is composed of three modelling choices and two algorithmic choices. The modelling choices are: (M.1) a model for the measurements, (M.2) an auxiliary process to model non-stationarity, and (M.3) a prior over model parameters conditioned on the auxiliary process and the past data. The algorithmic choices are: (A.1) an algorithm to compute an approximate posterior over the model parameters given the auxiliary variables, and (A.2) an algorithm to compute an approximate posterior—or more generally, a set of weights—over the auxiliary variables. We show how these different axes of variation span a wide variety of existing and new methods.111 BONE is designed for sequential prediction with time-indexed data, thus, it does not aim to find a “global” fit over all possible data points. Instead, BONE focuses on adaptive strategies that forecast the next datapoint (one at a time). This is in contrast to global optimisation approaches, such as Luo et al. (2024), which focus on finding the best global model for all datapoints. There are a number of probabilistic methods that handle non-stationarity and are not included in BONE. For example, Scalzo et al. (2021) and de Vilmarest & Wintenberger (2021) do not decouple the modelling of non-stationarity and the modelling over model parameters, thus, they cannot be written within our formulation. To illustrate this, Figure 1 shows how various methods can be written within the BONE framework. Here, we categorised methods by how the auxiliary variable is used to model non-stationarity and how it influences the prior over model parameters before observing a new data point.
Examples demonstrating the use of an easy-to-use library that implements these methods, written in Jax (Bradbury et al., 2018), are available at https://github.com/gerdm/BONE.
@@
[…]
Figure 1: Overview of BONE methods grouped by the qualitative nature of the auxiliary variable and the conditional prior. See Table 3 for a detailed breakdown of these methods.
@@ 2 The framework
Consider a sequence of measurements ${\bm{y}}_{1:T}=({\bm{y}}_{1},\ldots,{\bm{y}}_{T})$ with ${\bm{y}}_{i}\in^{d}$, and (unmodelled) features ${\bm{x}}_{1:T}=({\bm{x}}_{1},\ldots,{\bm{x}}_{T})$ with ${\bm{x}}_{i}\in^{q}$. Let $\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$ be a datapoint and $\mathcal{D}_{1:T}=(\mathcal{D}_{1},\ldots,\mathcal{D}_{T})$ the dataset at time $T$. Consider a conditional probabilistic model for ${\bm{y}}_{t}$, conditioned on ${\bm{x}}_{t}$, and parametrised by ${\bm{\theta}}_{t}\in^{m}$, given by $p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})$, and sometimes called the likelihood.
Assume that the conditional mean for ${\bm{y}}_{t}$ is encoded in a parametric model $h({\bm{\theta}}_{t},\,{\bm{x}}_{t})$, e.g., a neural network. A popular modelling choice for regression problems is to assume that $p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})={\cal N}({\bm{y}}_{t}\,|\,h({\bm{\theta}}_{t},\,{\bm{x}}_{t}),\mathbf{R}_{t})$ for known covariance $\mathbf{R}_{t}$. Next, assume there exists an auxiliary variable $\psi_{t}\in\boldsymbol{\Psi}_{t}$ that evolves following the dynamics $p(\psi_{t}\,|\,\psi_{t-1})$ and encodes information about the non-stationarity of the sequence at time $t$. Here, $\boldsymbol{\Psi}_{t}$ is the set of possible values of the auxiliary variable $\psi_{t}$. The purpose of this variable is, for instance, to determine which past datapoints ${\bm{y}}_{1:t-1}$ most closely align with the most recent measurement ${\bm{y}}_{t}$. We describe the auxiliary variable in detail in Section 2.4. Finally, the model parameters ${\bm{\theta}}_{t}$ evolve following the dynamics $p({\bm{\theta}}_{t}\,|\,{\bm{\theta}}_{t-1},\psi_{t})$. This represents how much parameters change, given the state of the auxiliary variable.
Figure 2 shows the probabilistic graphical model that motivates our formulation; this resembles the one in Doucet et al. (2000) with an additional optional dependence between the auxiliary variable and the measurements; in what follows we omit this dependence for brevity.
${\bm{\theta}}_{t-1}$ ${\bm{\theta}}_{t}$ ${\bm{\theta}}_{t+1}$ $\psi_{t-1}$ $\psi_{t}$ $\psi_{t+1}$ ${\bm{y}}_{t-1}$ ${\bm{y}}_{t}$ ${\bm{y}}_{t+1}$ ${\bm{x}}_{t-1}$ ${\bm{x}}_{t}$ ${\bm{x}}_{t+1}$
Figure 2: Two-levelled hierarchical state-space model (SSM) with known dynamics, motivating our BONE framework Solid arrows indicate required dependencies, while dashed arrows represent optional dependencies. Rectangles denote exogenous variables, and circles represent random variables. Observed elements are shaded in gray. The left shift in ${\bm{x}}_{t}$ represents that features are observed before observing ${\bm{y}}_{t}$.
For an experiment of length $T\in\mathbb{N}$, the joint conditional density over the model parameters, induced by the graphical model shown in Figure 2, is given by
$p({\bm{y}}_{1:T},{\bm{\theta}}_{0:T},\psi_{0:T}\,|\,{\bm{x}}_{1:T})=p({\bm{\theta}}_{0})\,p(\psi_{0})\,\prod_{t=1}^{T}p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})\,p({\bm{\theta}}_{t}\,|\,{\bm{\theta}}_{t-1},\psi_{t})\,p(\psi_{t}\,|\,\psi_{t-1}).$ (1)
@@
In this paper, we are interested in methods that efficiently compute the so-called expected posterior predictive $\hat{{\bm{y}}}_{t+1}:=\mathbb{E}_{p({\bm{\theta}}_{t},\psi_{t}\,|\,\mathcal{D}_{1:t})}[h({\bm{\theta}}_{t},{\bm{x}}_{t+1})]$ in an online and recursive manner. In our setting, one observes ${\bm{x}}_{t+1}$ just before observing ${\bm{y}}_{t+1}$; thus, to make a prediction about ${\bm{y}}_{t+1}$, we have ${\bm{x}}_{t+1}$ and $\mathcal{D}_{1:t}$ at our disposal.222 The input features ${\bm{x}}_{t+1}$ and output measurements ${\bm{y}}_{t+1}$ can correspond to different time steps. For example, ${\bm{x}}_{t+1}$ can be the state of the stock market at a fixed date and ${\bm{y}}_{t+1}$ is the return on a stock some days into the future. For the case of a discrete auxiliary variable $\psi_{t}\in\Psi_{t}$, the form of the expected posterior predictive for ${\bm{y}}_{t+1}$, induced by (1), is
$\hat{{\bm{y}}}_{t+1}$ $=\mathbb{E}_{p({\bm{\theta}}_{t},\psi_{t}\,|\,\mathcal{D}_{1:t})}[h({\bm{\theta}}_{t},{\bm{x}}_{t+1})]$ (2)
$=\sum_{\psi_{t}\in\Psi_{t}}\int h({\bm{\theta}}_{t},{\bm{x}}_{t+1})\,p({\bm{\theta}}_{t},\psi_{t}\,|\,\mathcal{D}_{1:t})\mathrm{d}{\bm{\theta}}_{t}$
$=\sum_{\psi_{t}\in\Psi_{t}}p(\psi_{t}\,|\,\mathcal{D}_{1:t})\int h({\bm{\theta}}_{t},{\bm{x}}_{t+1})p({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t})\mathrm{d}{\bm{\theta}}_{t},$
where
$p({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t})$ $\propto p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})\,p({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t-1}),$ (3)
$p({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t-1})$ $=\int p({\bm{\theta}}_{t}\,|\,{\bm{\theta}}_{t-1},\psi_{t})\,p({\bm{\theta}}_{t-1}\,|\,\mathcal{D}_{1:t-1})\,\mathrm{d}{\bm{\theta}}_{t-1},$ (4)
$p(\psi_{t}\,|\,\mathcal{D}_{1:t})$ $=p({\bm{y}}_{t}\,|\,{\bm{x}}_{t},\psi_{t},\mathcal{D}_{1:t-1})\,\sum_{\psi_{t-1}\in\boldsymbol{\Psi}_{t-1}}p(\psi_{t-1}\,|\,\mathcal{D}_{1:t-1})\,p(\psi_{t}\,|\,\psi_{t-1},\mathcal{D}_{1:t-1}).$ (5)
From (2), (3), (4), and (5) we argue that there are three key modelling choices and two algorithmic choices. Specifically, the three key modelling choices are: (M.1) the conditional mean $h({\bm{\theta}},{\bm{x}})$ together with the likelihood $p({\bm{y}}\,|\,{\bm{\theta}},{\bm{x}})$; (M.2) the auxiliary variable $\psi_{t}$; and (M.3) the conditional prior $p({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t-1})$. Additionally, the two algorithmic choices are: (A.1) the algorithm to compute (or approximate) the conditional posterior over model parameters $p({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t})$, and (A.2) the algorithm that computes (or approximates) the posterior over weights $p(\psi_{t},\,|\,\mathcal{D}_{1:t})$.
The BONE framework generalises these choices, allowing for greater flexibility while maintaining the motivating probabilistic structure. Instead of the likelihood model $p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})$ with conditional mean $h({\bm{\theta}}_{t},{\bm{x}}_{t})$, we consider a general function $\exp(-\ell({\bm{y}}_{t};\,{\bm{\theta}}_{t},{\bm{x}}_{t}))$, where $\ell({\bm{y}}_{t};\,{\bm{\theta}}_{t},{\bm{x}}_{t})$ could be either a loss function or a log-likelihood. Next, instead of the conditional prior $p({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t-1})$, we introduce a more general modelling function, $\pi({\bm{\theta}}_{t};\psi_{t},\mathcal{D}_{1:t-1})$ that governs the prior over model parameters.333 This function adopts an ad hoc approach to parameter evolution instead of explicitly solving the integration step (4). In Appendix C, we detail how the inference is carried out when the dynamics are known. More precisely, we conduct a full-Bayesian treatment by calculating the posterior of model parameters at $t+1$ and using this for inference. Similarly, instead of the posterior density $p({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t})$, we employ the function $q({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})$; e.g., an approximation of the posterior, or a generalised posterior (Bissiri et al., 2016). Finally, instead of the posterior over weights $p(\psi_{t}\,|\,\mathcal{D}_{1:t})$, we consider a weighting function $\nu(\psi_{t};\,\mathcal{D}_{1:t})$, which can be the Bayesian posterior or an ad-hoc time-dependent weighting function.
This generalisation is important because it unifies a wide range of existing methods under a common framework. Many well-known approaches in the literature can be written as elements of BONE by appropriately selecting the model for measurements, the conditional prior, and the posterior approximations. BONE highlights connections between different methods, it also enables systematic comparisons under a common umbrella, and it allows us to develop novel algorithms. Table 1 explicitly contrasts the choices in BONE with those in the classical Bayesian formalism.
component BONE Bayes
(M.1: model) $h({\bm{\theta}}_{t},{\bm{x}}_{t})$  &  $\exp(-\ell({\bm{y}}_{t};{\bm{\theta}}_{t},{\bm{x}}_{t}))$ $h({\bm{\theta}}_{t},{\bm{x}}_{t})$  &  $p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})$
(M.2: auxvar) $\psi_{t}$ $\psi_{t}$
(M.3: prior) $\pi_{t}({\bm{\theta}}_{t};\,\psi_{t}):=\pi({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t-1})$ $p({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t-1})$
(A.1: posterior) $q_{t}({\bm{\theta}}_{t};\psi_{t}):=q({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})$ $p({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t})$
(A.2: weighting) $\nu_{t}(\psi_{t}):=\nu(\psi_{t};\,\mathcal{D}_{1:t})$ $p(\psi_{t}\,|\,\mathcal{D}_{1:t})$
Table 1: Components of the BONE framework.
With these modifications, the expected posterior predictive under BONE is
$\hat{{\bm{y}}}_{t}$ $:=\sum_{\psi_{t}\in\Psi_{t}}\underbrace{\nu(\psi_{t}\,|\,\mathcal{D}_{1:t})}_{\text{(A.2: weighting)}}\,\int\underbrace{h({\bm{\theta}}_{t},{\bm{x}}_{t+1})}_{\text{(M.1: model)}}\,\underbrace{q({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})}_{\text{(A.1: posterior)}}\mathrm{d}{\bm{\theta}}_{t},$ (6)
where
$q({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})\propto\underbrace{\pi({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t-1})}_{\text{(M.3: prior)}}\,\underbrace{\exp(-\ell({\bm{y}}_{t};{\bm{\theta}}_{t},{\bm{x}}_{t}))}_{\text{(M.1: model)}}$ (7)
takes the form of a generalised posterior (Bissiri et al., 2016). In classical Bayesian setting, the loss function takes the form of the negative log-likelihood, i.e.,
$\ell({\bm{y}}_{t};{\bm{\theta}}_{t},{\bm{x}}_{t})=-\log p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t}).$ (8)
Unless stated otherwise, we work with the negative log-likelihood in (8) found in the classical Bayesian setting.
A prediction for ${\bm{y}}_{t+1}$ given $\mathcal{D}_{1:t}$, ${\bm{x}}_{t+1}$, and $\psi_{t}$ is
$\hat{{\bm{y}}}_{t+1}^{(\psi_{t})}=\mathbb{E}_{q_{t}}[h({\bm{\theta}}_{t};\,{\bm{x}}_{t+1})\,|\,\psi_{t}]:=\int h({\bm{\theta}}_{t};\,{\bm{x}}_{t+1})\,q({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})\mathrm{d}{\bm{\theta}}_{t}\,.$ (9)
@@
Here we use the shorthand notation $q_{t}=q({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})$ and $\mathbb{E}_{q_{t}}[\cdot\,|\,\psi_{t}]$ to highlight dependence on $\psi_{t}$.
@@
Algorithm 1 provides pseudocode for the prediction and update steps in the BONE framework. Notably, these components can be broadly divided into two categories: modelling and algorithmic. The modelling components determine the inductive biases in the model, and correspond to $h$, $\ell$, $\psi_{t}$, and $\pi$. The algorithmic components dictate how operations are carried out to produce a final prediction — this corresponds to $q_{t}$ and $\nu_{t}$.
@@
0:  $h({\bm{\theta}},{\bm{x}}_{t})$ // Choice of (M.1: model)
@@
2:     $\pi_{t}({\bm{\theta}}_{t};\,\psi_{t})\leftarrow\pi({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t-1})$ // choice of (M.3: prior)
3:     $q_{t}({\bm{\theta}}_{t};\,\psi_{t})\leftarrow q({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})\propto\pi_{t}({\bm{\theta}}_{t}\;\psi_{t})\,\exp(-\ell({\bm{y}}_{t};{\bm{\theta}}_{t},{\bm{x}}_{t}))$// choice of (A.1: posterior)
4:     $\nu_{t}(\psi_{t})\leftarrow\nu(\psi_{t};\,\mathcal{D}_{1:t})$ // choice of (A.2: weighting)
5:     $\hat{{\bm{y}}}_{t+1}^{(\psi_{t})}\leftarrow\mathbb{E}_{q_{t}}[h({\bm{\theta}}_{t},{\bm{x}}_{t+1});\,\psi_{t}]$ // conditional prequential prediction